OpenStudy (anonymous):

Suppose that the graph of f contains the point (4,5). Find a point that must be on the graph g(x)=2f(x+6).

OpenStudy (anonymous):

if \(x+6=4\) then \(x=-2\)

OpenStudy (anonymous):

so if \(x=-2\) then \(2(f(-2+6)=2f(4)=2\times 5=10\) meaning \((-2,10)\) is on the graph of \(g(x)=2f(x+6)\)

OpenStudy (anonymous):

thank you!!!